State the phase shift of y=cos(theta-pi/3). Then graph the function. The graph that it asks me to use has pi and 2pi as the x coordinates, I'm not sure how to graph on it.

Answer by lwsshak3(11628)

State the phase shift of y=cos(theta-pi/3). Then graph the function. The graph that it asks me to use has pi and 2pi as the x coordinates, I'm not sure how to graph on it.
***
Form of equation for cos: y=Acos(Bx-C), A=amplitude, period=2π/B, phase shift=(C/B)
For given equation: y=cos(x-π/3)
amplitude=1
B=1
period=2π/B=2π
(1/4) period=π/2
phase shift: C/B=π/3/1=π/3 (shift to the right)
Graphing:
start with coordinates of basic cos curve for one period: (x-axis scale in radians)
(0,1), ((π/2),0), (π,-1), (3π/2,(-1)), (2π,0)
shift curve (π/3) to the right
(0,.5), ((π/3),1), ((5π/6),0), (4π3,-1), (11π/6,(-1)), (7π/3,0)